# MSP-MVS: Multi-Granularity Segmentation Prior Guided Multi-View Stereo

### v5

@@ Introduction
Multi-View Stereo (MVS) is a technique that employs images captured from different viewpoints to reconstruct 3D scenes or objects. Its diverse application and downstream task spans across object detection (Lu et al. 2023, 2024), image denoising (Yang et al. 2024; Li et al. 2022), video representation (Liu et al. 2024a, b), multiview clustering (Guan et al. 2024b, a), etc. Numerous ideas (Chen et al. 2024a, b; Kang et al. 2025) and datasets (Yao et al. 2020) have advanced its progress. However, it still faces the primary challenge of reconstructing large textureless areas.
@@
Recently, the advancement of deep learning has led to the emergence of numerous large-language models (Luo et al. 2024a, b), gaussian splatting algorithms (Gan et al. 2023a, b), and self-supervised learning methods (Li and Zhang 2022; Song et al. 2025; Chen et al. 2023, 2024c; Cheng and Sun 2024; Sun et al. ) and vision-language model (Chen et al. 2024d; Qian et al. 2024; Chen et al. 2024e). In MVS field, MVSNet (Yao et al. 2018) is the first to utilize differentiable 3D cost volumes to construct deep neural networks. To reduce memory, Cas-MVSNet (Yang et al. 2020) adopts a coarse-to-fine strategy to retain depths across multiple scales. IterMVS-LS (Wang et al. 2022a) leverages a lightweight probability estimator to encode depth distribution during regularization. For feature extraction, AAR-MVSNet (Wei et al. 2021) employed deformable convolutions to achieve adaptive feature aggregation. Differently, MVSTER (Wang et al. 2022b) leverages the transformer structure to introduce the multi-head attention mechanism. Geo-MVSNet (Zhang et al. 2023b) adopts a two-branch geometry fusion network to enhance geometry perception. RA-MVSNet (Zhang et al. 2023a) associates hypothesis planes with surface patches to enhance perception field. Moreover, Raymvsnet++ (Shi et al. 2023) applies an epipolar transformer to enhance feature aggregation along camera rays. GoMVS (Wu et al. 2024) integrates a geometrically consistent propagation module to refine cost aggregation. Differently, our method targets on aggregating multi-granularity depth edges for edge-confined patch deformation. Despite this, most of them still suffer from either unaffordable memory costs or limited generalization.
@@
The conventional PM method which utilizes fixed patch for reconstruction should be detailed at first. Given a series of input images =$\mathcal{I}=\{I_{n}\mid n=1\cdots N\}$ and corresponding camera parameters $\mathcal{P}=\{K_{n},R_{n},C_{n}\mid n=1\cdots N\}$, each image is selected from $\mathcal{I}$ to be the reference image $I_{i}$, with its depth map sequentially reconstructed through pairwise matching with the remaining images $(\mathcal{I}-I_{i})$. Here $K$, $R$ and $C$ respectively represent the intrinsic parameters, the rotation matrix and the camera center. Specifically, given reference image $I_{i}$ and source image $I_{j}$, the homography matrix $H_{ij}$ for pixel $p$ in $I_{i}$ is defined as follows:
@@
where $n^{T}$ and $d$ respectively denote normal and depth hypothesis. Given the fixed patch which is a square window $W$ centered on $p$ in $I_{i}$, we can leverage $H_{ij}$ to obtain its corresponding patch $W^{\prime}$ in $I_{j}$ Then the matching cost is:
@@
where $cov$ is weighted covariance (Schönberger et al. 2016). Finally, we can obtain multi-view aggregated cost via view weights (Xu and Tao 2019) and further leverage propagation and refinement to update depth with the minimum cost.
@@
where $\mathbb{I}(\cdot)$ is the indicator function such that $\mathbb{I}(\text{true})=1$ and $\mathbb{I}(\text{false})=0$. Therefore, reliable masks that maximize the quantity of $S_{max}$ are effectively retained. Finally, as shown in Fig. 3 (i), by aggregating all label masks among all layers, we effectively obtain the aggregated image, which is then further used by function $\tau$ to derive the aggregated boundary map: $B^{\prime}=\tau(\bigcup_{k=1}^{K}\sum_{n=1}^{N}M_{n}^{k})$. $B^{\prime}$ is then adopted as the multi-granularity segmentation prior to constrain each unreliable pixel $p$ in $I_{i}$ to obtain the new anchor collection $S^{\prime}$ for patch deformation as shown in Fig. 3 (d). Although $S^{\prime}$ contains fewer anchors than $S$, it significantly enhances the anchor’s reliability. Compared with APD-MVS in Fig. 3 (a), our edge-confined patch deformation in Fig. 3 (d) effectively avoids unexpected edge-skipping and constraints patch deformation in homogeneous areas, thus avoiding distortion.
@@
Specifically, as shown in Fig. 4 (b), centered on unreliable pixel $p$, we first emit 8 rays at fixed angle $45^{\circ}$ for sector division. We then define the first reliable pixel or depth edge encountered by each ray as the boundary pixel. Thus, we have all sector angles $\Theta=\{\theta_{i}\mid i=1\cdots 8\}$, sectors $\mathcal{V}=\{V_{i}\mid i=1\cdots 8\}$, and boundary pixels =$\mathcal{B}=\{b_{i},b_{i+1}\mid i=1\cdots 8\}$. Subsequently, for sector $V_{i}$, if either its boundary pixels $b_{i}$ or $b_{i+1}$ is located on depth edges rather than reliable pixels, we consider $V_{i}$ to be fully enclosed by depth edges and contain no reliable pixels. Otherwise, we consider $V_{i}$ to be fully enclosed by a certain number of reliable pixels, denoted as $n_{i}$. Since reliable pixels are typically continuously distributed, $n_{i}$ can be approximately estimated as the L1 distance between two boundary pixels $b_{i}$ and $b_{i+1}$. Moreover, we can obtain the average number of all reliable pixels $\bar{d}$ among all sectors $\mathcal{V}$. Then through the weighted averaging of all reliable pixels, we can proportionally recalculate all angles $\Theta^{\prime}=\{\theta_{i}^{\prime}\mid i=1\cdots 8\}$ and in turn acquire new sectors $\mathcal{V}^{\prime}$ and boundary pixels $\mathcal{P}^{\prime}$ for allocation. The $k^{th}$ rectified angle $\theta_{k}^{\prime}$ is calculated by:

### v6

@@ Introduction
Multi-View Stereo (MVS) is a technique that employs images captured from different viewpoints to reconstruct 3D scenes or objects. Its diverse application and downstream task spans across object detection (Lu et al. 2023, 2024), image denoising (Yang et al. 2024c; Li et al. 2022a), video representation (Liu et al. 2024b, c), multiview clustering (Guan et al. 2024b, a), etc. Numerous ideas (Chen et al. 2024a, b; Kang et al. 2025) and datasets (Yao et al. 2020) have advanced its progress. However, it still faces the primary challenge of reconstructing large textureless areas.
@@
Recently, the advancement of deep learning has led to the emergence of numerous large-language models (Luo et al. 2024b, c), gaussian splatting algorithms (Gan et al. 2023a, b), and self-supervised learning methods (Li and Zhang 2022; Song et al. 2025c; Chen et al. 2023, 2024c; Cheng and Sun 2024; Sun et al. 2024) and vision-language model (Chen et al. 2024d; Qian et al. 2024; Chen et al. 2024e). In MVS field, MVSNet (Yao et al. 2018) is the first to utilize differentiable 3D cost volumes to construct deep neural networks. To reduce memory, Cas-MVSNet (Yang et al. 2020) adopts a coarse-to-fine strategy to retain depths across multiple scales. IterMVS-LS (Wang et al. 2022a) leverages a lightweight probability estimator to encode depth distribution during regularization. For feature extraction, AAR-MVSNet (Wei et al. 2021) employed deformable convolutions to achieve adaptive feature aggregation. Differently, MVSTER (Wang et al. 2022b) leverages the transformer structure to introduce the multi-head attention mechanism. Geo-MVSNet (Zhang et al. 2023c) adopts a two-branch geometry fusion network to enhance geometry perception. RA-MVSNet (Zhang et al. 2023b) associates hypothesis planes with surface patches to enhance perception field. Moreover, Raymvsnet++ (Shi et al. 2023) applies an epipolar transformer to enhance feature aggregation along camera rays. GoMVS (Wu et al. 2024) integrates a geometrically consistent propagation module to refine cost aggregation. Differently, our method targets on aggregating multi-granularity depth edges for edge-confined patch deformation. Despite this, most of them still suffer from either unaffordable memory costs or limited generalization.
@@
The conventional PM method which utilizes fixed patch for reconstruction should be detailed at first. Given a series of input images $\mathcal{I}=\{I_{n}\mid n=1\cdots N\}$ and corresponding camera parameters $\mathcal{P}=\{K_{n},R_{n},C_{n}\mid n=1\cdots N\}$, each image is selected from $\mathcal{I}$ to be the reference image $I_{i}$, with its depth map sequentially reconstructed through pairwise matching with the remaining images $(\mathcal{I}-I_{i})$. Here $K$, $R$ and $C$ respectively represent the intrinsic parameters, the rotation matrix and the camera center. Specifically, given reference image $I_{i}$ and source image $I_{j}$, the homography matrix $H_{ij}$ for pixel $p$ in $I_{i}$ is defined as follows:
@@
where $n^{T}$ and $d$ respectively denote normal and depth hypothesis. Given the fixed patch which is a square window $W$ centered on $p$ in $I_{i}$, we can leverage $H_{ij}$ to obtain its corresponding patch $W^{\prime}$ in $I_{j}$ . Then the matching cost is:
@@
where covcov is weighted covariance (Schönberger et al. 2016). Finally, we can obtain multi-view aggregated cost via view weights (Xu and Tao 2019) and further leverage propagation and refinement to update depth with the minimum cost.
@@
where $\mathbb{I}(\cdot)$ is the indicator function such that $\mathbb{I}(\text{true})=1$ and $\mathbb{I}(\text{false})=0$. Therefore, reliable masks that maximize the quantity of $S_{max}$ are effectively retained. Finally, as shown in Fig. 3 (i), by aggregating all label masks among all layers, we effectively obtain the aggregated image, which is then further used by function $\tau$ to derive the aggregated boundary map: $^{\prime}=\tau(\bigcup_{k=1}^{K}\sum_{n=1}^{N}M_{n}^{k}). BB^{\prime}$ is then adopted as the multi-granularity segmentation prior to constrain each unreliable pixel $p$ in $I_{i}$ to obtain the new anchor collection $S^{\prime}$ for patch deformation as shown in Fig. 3 (d). Although $S^{\prime}$ contains fewer anchors than $S$, it significantly enhances the anchor’s reliability. Compared with APD-MVS in Fig. 3 (a), our edge-confined patch deformation in Fig. 3 (d) effectively avoids unexpected edge-skipping and constraints patch deformation in homogeneous areas, thus avoiding distortion.
@@
Specifically, as shown in Fig. 4 (b), centered on unreliable pixel $p$, we first emit 8 rays at fixed angle $45^{\circ}$ for sector division. We then define the first reliable pixel or depth edge encountered by each ray as the boundary pixel. Thus, we have all sector angles $\Theta=\{\theta_{i}\mid i=1\cdots 8\}$, sectors $\mathcal{V}=\{V_{i}\mid i=1\cdots 8\}$, and boundary pixels $\mathcal{B}=\{b_{i},b_{i+1}\mid i=1\cdots 8\}$. Subsequently, for sector $V_{i}$, if either its boundary pixels $b_{i}$ or $b_{i+1}$ is located on depth edges rather than reliable pixels, we consider $V_{i}$ to be fully enclosed by depth edges and contain no reliable pixels. Otherwise, we consider $V_{i}$ to be fully enclosed by a certain number of reliable pixels, denoted as $n_{i}$. Since reliable pixels are typically continuously distributed, $n_{i}$ can be approximately estimated as the L1 distance between two boundary pixels $b_{i}$ and $b_{i+1}$. Moreover, we can obtain the average number of all reliable pixels $\bar{d}$ among all sectors $\mathcal{V}$. Then through the weighted averaging of all reliable pixels, we can proportionally recalculate all angles $\Theta^{\prime}=\{\theta_{i}^{\prime}\mid i=1\cdots 8\}$ and in turn acquire new sectors $\mathcal{V}^{\prime}$ and boundary pixels $\mathcal{P}^{\prime}$ for allocation. The $k^{th}$ rectified angle $\theta_{k}^{\prime}$ is calculated by:
@@ Acknowledgements
Supplementary Material
In this supplementary material, we significantly expand upon the related work discussed in the main paper to provide a comprehensive view of the research landscape surrounding MSP-MVS. We structure this discussion into two main sections. The first section, ”Expanded Related Work,” situates our method within the direct lineage of 3D computer vision, from classical geometric principles to modern neural representations. The second section, ”Broader Context and Future Directions,” elevates the perspective, connecting MSP-MVS to the paradigm-shifting developments in foundation models and cross-disciplinary AI research, thereby outlining a vision for the future of 3D perception.
More Details on Implementation
Results on ETH3D and TNT dataset
Fig. 8 demonstrates qualitative comparisons among various methods on partial scenarios of the ETH3D datasets. Similarly, Fig. 9 and Fig. 10 illustrate the qualitative results in the intermediate and advanced sets of the Tanks & Temples datasets, respectively. As can be seen, our method reconstructs the most complete and realistic point clouds than other state-of-the-art traditional MVS methods, especially in recovering large textureless areas such as walls and floors, without introducing detail distortion, evidencing its strong effectiveness and robust generalization ability.
Figure 8: An illustration of the qualitative results on partial scenes of ETH3D datasets (terrace 2, lounge, pipes, meadow, lecture room, exhibition hall, living room, relief and relief 2). It is obvious that our methods outperform others, especially in large textureless areas.
Figure 9: Reconstructed point clouds results on intermediate set of TNT datasets.
Figure 10: Reconstructed point clouds results on advanced set of TNT datasets.
Analyses and Discussions
Q1: Compared to fixed interval sampling or global sampling, what is the success rate or accuracy of the adaptive sampling strategy in Synergistic 3D Optimization?
A1: As shown in Tab. 4, we respectively perform the fixed interval sampling strategy (Fix.), the global sampling strategy (Glo.) and the adaptive sampling strategy (Apt.) on deformed patches of ETH3D training dataset for comparison. Evidently, the F1 scores of Apt. significantly outperform Fix., indicating the effectiveness of adaptive sampling strategy. While Glo. achieves a slightly better F1 score than Apt., it requires over 5 times the runtime compared to our method, indicating that our adaptive sampling strategy can achieve better trade-offs between time efficiency and performance.
Moreover, as shown in Fig. 11, we present different 2D cost profiles of deformed patches employed by different sampling strategies for intuitive comparison. Obviously, the fixed interval sampling strategy (i.e. green line) inevitably causes cost to fall into a local-minimum. Additionally, while the cost profile of global sampling strategy (i.e. blue line) achieves a lower overall cost, both the global and adaptive sampling strategies (i.e. red line) can approximate the global-minimum matching cost. Consequently, our adaptive sampling strategies can identify the global minimum cost with less time, demonstrating its high efficiency.
Q2: There are already MVS methods based on semantic segmentation like SD-MVS (Yuan et al. 2024a) and Multi-View SAM (Shvets et al. 2024), what’s our advantage?
A2: While existing methods like SD-MVS and Multi-View SAM utilize semantic segmentation, they typically leverage semantic information at a single granularity, which often fails to capture comprehensive depth edges in complex scenarios. In contrast, our approach can integrate and further refine semantic information across multiple granularities, allowing us to aggregate more robust depth edges. For instance, when reconstructing a book, text and images on its cover might be mistakenly identified as depth edges by SD-MVS. However, our MSP-MVS can effectively exclude such distractions, treating them together as homogeneous areas for patch deformation, thus improving matching accuracy.
Q3: In the Disassemble-Clustering strategy, what function does the disassemble operation serve?
A3: The disassemble operation in disassemble-clustering strategy serves to maintain the overall patch size while allowing our method to select more reliable anchors. By quadrupling the number of anchors and disassembling the original patch into multiple smaller sub-patches, the disassemble operation guarantees the deformed patches covering more reliable pixels without increasing computational complexity.
Method $2cm$ $10cm$
F1 Comp. Acc. F1 Comp. Acc.
Fix. 86.94 87.29 86.8 96.9 97.33 96.47
Glo. 88.45 88.32 88.78 97.94 98.03 97.85
Apt. 88.09 87.87 88.51 97.70 97.74 97.67
Table 4: Quantitative ablation studies on ETH3D dataset of different sampling strategies employed in deformed patches.
Q4: How exactly does the conditional random field algorithm achieve edge refinement?
A4: According to (Lafferty, McCallum, and Pereira 2001), the conditional random field can leverage inter-pixel correlation to refine label distribution, which is achieved by modeling the predictions within a graphical model that illustrates the dependencies among them.
Specifically, in the main text, we have already defined the unary potential $\psi_{u}\left(l_{i}\right)$ and the pairwise potential $\psi_{p}\left(l_{i},l_{j}\right)$. We then obtain the overall energy function $\Psi(L)$ by aggregating $\psi_{u}\left(l_{i}\right)$ and $\psi_{p}\left(l_{i},l_{j}\right)$ among all pixels, calculated by:
$\Psi(L)=\sum_{i}^{K}\psi_{u}\left(l_{i}\right)+\sum_{i,j}^{K}\psi_{p}\left(l_{i},l_{j}\right).$ (10)
Figure 11: Comparative 2D cost profiles of different sampling strategies. The green, blue, red line in left respectively denote fixed interval, global and our adaptive sampling strategy employed in red deformable PM anchors of right graph.
By minimizing $\Psi(L)$ CRF effectively achieves edge refinement. To accomplish the minimization process, an observation sequence $X$ and a corresponding label sequence $Y$ are defined to express the conditional probability of the model:
$P(Y\mid X)=\frac{1}{Z(X)}\exp\left(\sum_{k}\lambda_{k}\Psi_{k}\left(Y,X\right)\right),$ (11)
where $X=\{x_{1},x_{2},...,x_{n}\}$ and $Y=\{y_{1},y_{2},...,y_{n}\}$ respectively represents all pixels and their corresponding labels, $\Psi_{k}\left(Y,X\right)$ denotes the feature function, each feature function corresponds to a weight $\lambda_{k}$, and $Z(X)$ is the normalization factor to ensure a valid probability distribution, computed by summing over all possible label configurations $Y$. Subsequently, to maximize the prediction accuracy $\mathcal{L}(\Lambda)$, the algorithm employs the maximization of the log-likelihood function to optimize the parameters $\lambda_{k}$. Mathematically, $\mathcal{L}(\Lambda)$ is calculated by:
$\mathcal{L}(\Lambda)=\sum_{(X,Y)}\log P(Y\mid X;\Lambda)-\frac{1}{2\eta^{2}}\|\Lambda\|^{2},$ (12)
where $\Lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{k}\}$ represents the set of all feature function weights, $\eta$ is a regularization term to prevent overfitting. The maximization of $\mathcal{L}(\Lambda)$ can be achieved through the Quasi-Newton method (Sutton and McCallum 2012).
Expanded Related Work
This section contextualizes MSP-MVS within the broader evolution of Multi-View Stereo (MVS) and 3D scene representation. We begin by delving into the advancements in patch-based MVS, transition to the rise of learning-based 3D representations, discuss the critical role of semantic and geometric priors, and conclude with the diverse applications of these technologies, illustrating the impact and relevance of robust 3D reconstruction.
Advancements in Patch-based and Hybrid MVS
Traditional MVS methods have long relied on the PatchMatch algorithm (Barnes et al. 2009), which iteratively refines depth estimates by propagating and matching local patches. However, these methods often falter in texture-poor regions where patch-based matching becomes ambiguous. To address this, hybrid approaches have emerged, augmenting geometric principles with more robust strategies. Our work, MSP-MVS, falls into this category by introducing multi-granularity segmentation priors to guide patch deformation. This line of research, focusing on enhancing reconstruction in challenging areas, has seen significant recent activity. For instance, similar to our use of semantic edges, other works have proposed using dual-level precision edges to guide MVS and ensure accurate planarization (Chen et al. 2025a). The core idea of leveraging segmentation for robust reconstruction is further explored in methods that are explicitly segmentation-driven, incorporating sophisticated mechanisms like depth restoration and occlusion handling to improve results (Yuan et al. 2025c). An earlier precursor to these ideas can be seen in textureless-aware segmentation and correlative refinement guided MVS, which also highlights the value of semantic understanding in traditional MVS pipelines. These methods collectively demonstrate a clear trend: the fusion of high-level semantic guidance with low-level geometric matching is crucial for achieving state-of-the-art performance.
The Rise of Learning-based 3D Representations
While hybrid methods enhance traditional pipelines, the field has been revolutionized by end-to-end learning-based approaches. This paradigm shift began with MVS networks that replaced handcrafted cost functions with learned 3D cost volumes. Recent advancements in this domain have focused on improving network architectures, for instance, by leveraging monocular priors to guide the multi-view process (Jiang et al. 2025b) or by incorporating powerful backbones like Recurrent Regularization Transformers (Jiang et al. 2025c) and State Space Models (Jiang et al. 2025a) to better aggregate information across views.
Concurrently, a new frontier opened with the development of implicit and explicit neural representations for scenes. Neural Radiance Fields (NeRF) offered a way to synthesize photorealistic novel views from a continuous volumetric function, and community efforts have produced comprehensive datasets to benchmark such NeRF-based 3D reconstruction techniques (Yan et al. 2023). More recently, 3D Gaussian Splatting (GS) has gained immense popularity due to its ability to achieve real-time rendering with high fidelity. Research in this direction is rapidly expanding, with a focus on several key areas. These include developing lightweight and compact models for large-scale dynamic 4D scenes through context-aware modeling (Liu et al. 2025c, c), enabling real-time video representation at high frame rates through deformable 2D Gaussians (Liu et al. 2025d), and creating intuitive tools for controllable 3D scene editing (Yan et al. 2024b). Furthermore, efforts are being made to enhance the structural integrity of GS reconstructions by leveraging topological concepts like persistent homology (Shen et al. 2025), and to unify appearance codes with bilateral grids for complex driving scene reconstruction (Wang et al. 2025d). A critical challenge is extending these representations to encompass not just static scenes but the world’s dynamics, leading to the development of world models with self-supervised 3D labels for robotics applications (Yan et al. 2025). Bridging the gap between these new representations and classic computer vision tasks, recent work has also focused on performing object segmentation directly on the Gaussian representation, guided by gradient information to enhance boundary precision (Li et al. 2024), and enabling spatio-temporal decoupling for real-time dynamic scene rendering (Li et al. 2025d).
Semantic and Geometric Priors in 3D Perception
The core innovation of MSP-MVS is the use of multi-granularity priors from Semantic-SAM. This highlights a broader principle: leveraging structured information is key to solving ill-posed perception problems. This principle is widely applied across various domains. In the context of segmentation, for instance, researchers have explored scaling up autoregressive pretraining for complex structures like neurons (Chen et al. 2025e), using dual-form complementary masking for domain-adaptive image segmentation (Wang et al. 2025a), and employing multi-agent reinforcement learning for self-supervised segmentation (Chen et al. 2023).
Beyond segmentation, graph-based models offer a powerful way to represent structural relationships. Techniques such as multi-scale graph learning for challenging tasks like anti-sparse downscaling (Fan et al. 2025) demonstrate the versatility of graph structures. In the realm of graph autoencoders, revisiting masking strategies from a robustness perspective (Song et al. 2025b) and developing self-purified models to release robust expression power (Song et al. 2025d) have pushed the boundaries of representation learning. For composed image and video retrieval, a deep understanding of fine-grained semantics is paramount. This has led to methods that explicitly parse modification semantics (Li et al. 2025c), perform visual-semantic decomposition and partial alignment (Qu et al. 2024), disentangle representations with complementarity guidance (Fu et al. 2025), utilize hierarchical uncertainty-aware disambiguation (Chen et al. 2025g), revise focus shifts based on segmentation (Chen et al. 2025f), and mine entity and modification relations (Li et al. 2025b). These diverse approaches underscore the universal importance of leveraging semantic and structural priors, a principle that lies at the heart of our MSP-MVS.
Applications of 3D Reconstruction and Perception
The ultimate goal of advancing 3D perception is to enable robust real-world applications. Autonomous Driving is a primary beneficiary. Accurate 3D reconstruction and perception are essential for building online HD maps (Zhang et al. 2025a), predicting trajectories from a cross-view perspective (Song et al. 2023), and estimating depth under adverse weather conditions through curriculum contrastive learning (Wang et al. 2024a). This same line of inquiry has also explored the synergy between contrastive learning and diffusion models for even more robust depth estimation, demonstrating the rapid evolution of perception techniques in this area (Wang et al. 2024b). Recent work even explores transforming image editors into dense geometry estimators (Wang et al. 2025c) and leveraging adversarial transfer for end-to-end driving models (Zhang et al. 2025c). Simplified yet effective models with distinct experts and implicit interactions are also being developed to tackle the complexity of end-to-end driving (Zhang et al. 2025d).
In Remote Sensing, multi-view graph clustering with dual relation optimization (Guan et al. 2025a), structure-adaptive multi-view graph clustering (Guan et al. 2025c), and sampling-enhanced contrastive clustering with long-short range information mining (Guan et al. 2025b) are all techniques developed to handle the unique challenges of satellite and aerial imagery. The development of interactive agents for comprehensive change interpretation and analysis (Liu et al. 2024a) and comprehensive surveys on spatiotemporal vision-language models for remote sensing (Liu et al. 2025b) further highlight the rapid progress in this field. Diffusion models are also being used for controllable image generation (Tang et al. 2024) and to enhance object detection through data generation (Tang et al. 2025).
Finally, in Robotics and Human-Object Interaction (HOI), discovering syntactic interaction clues (Luo et al. 2024a) and employing synergistic prompting learning (Luo et al. 2025a) are crucial for enabling machines to understand and interact with their environment in a human-like manner.
In summary, this expanded review situates MSP-MVS at the confluence of several critical research streams. By building upon the robust foundations of patch-based MVS while integrating high-level semantic priors, our work addresses long-standing challenges in textureless reconstruction. It stands as a complementary approach to the concurrent rise of learning-based representations like NeRF and Gaussian Splatting, offering a pathway that prioritizes geometric accuracy and computational efficiency. As demonstrated by the diverse and demanding applications in autonomous driving and remote sensing, the need for reliable and precise 3D perception remains paramount. MSP-MVS contributes a significant step forward in this direction, proving that the principles of guided geometric matching remain highly relevant in the modern 3D vision landscape.
Broader Context and Future Directions: The Role of Foundation Models and Cross-Disciplinary Insights
Having situated MSP-MVS within the context of 3D vision, we now broaden our perspective to discuss its connections to the wider trends in artificial intelligence. This section explores how the rise of foundation models, advances in general learning paradigms, and insights from specialized, cross-disciplinary domains are poised to shape the future of MVS and 3D perception at large.
The Era of Vision-Language Models (VLMs): Towards Language-Guided 3D Systems
The rise of Large Language Models (LLMs) and Vision-Language Models (VLMs) is reshaping computer vision. The ability of these models to reason, follow instructions, and process multi-modal information offers exciting possibilities for MVS. Our use of Semantic-SAM is a step in this direction, but future systems could be guided by more complex, natural language commands. Research into VLM capabilities is vast, spanning from enhancing LLMs’ confidence on edited facts (Bi et al. 2024b) and aligning them for context-faithfulness (Bi et al. 2024a) to controlling knowledge reliance through fine-grained mechanisms (Bi et al. 2025c) and promoting exploration to improve multi-domain reasoning (Bi et al. 2025b).
VLMs are being applied to a wide array of tasks that hint at their potential for 3D. This includes versatile advertising poster layout generation (Anonymous 2025), explainable visual question answering via diffusion-based chain-of-thought (Lu, Lu, and Luo 2024), and challenging VLMs’ spatial intelligence through complex reasoning tasks (Song et al. 2025e). For specific applications, VLMs are being used for document-based zero-shot learning by progressively optimizing prompts (Qu et al. 2025a) and leveraging multi-attribute document supervision (Qu et al. 2025b). Instruction-based learning is also proving effective for complex tasks like HOI detection (Luo et al. 2025b). The synergy between retrieval mechanisms and preference optimization is being used to better align VLMs (Xing et al. 2025), and hierarchical cross-modal alignment is being developed for decoupled multimodal representation learning (Qian et al. 2025c). The most direct link to 3D is seen in generative text-guided 3D vision-language pretraining for medical image segmentation (Chen et al. 2025d), which paves the way for future language-guided 3D reconstruction systems.
Advanced Learning Paradigms and Their Implications for MVS
Beyond VLMs, advances in general machine learning paradigms offer promising directions for MVS. Data-centric AI emphasizes the quality of training data. This includes learning to refine pre-training data at scale from expert-guided programs (Bi et al. 2025a) and developing robust digital watermarking frameworks resistant to extreme cropping, scaling (Sun et al. 2025b), and non-differentiable distortions (Sun et al. 2025a) to ensure data integrity.
In terms of model architectures and training, new models like the diffusion Mamba for end-to-end multi-modal tasks (Lu et al. 2025) and lightweight model transition methods to improve LLM training via weight scaling (Li et al. 2025a) could lead to more efficient and powerful MVS networks. Furthermore, in the realm of graph-based learning, research into elevating the resilience of graph prompt tuning against adversarial attacks (Song et al. 2025a) is critical for building robust systems.
Finally, the concept of intelligent agents is gaining traction. This includes the development of world cognition agents enabling adaptive human-AI symbiosis in industrial settings (Liu et al. 2025a) and comprehensive surveys on pure Vision Language Action (VLA) models (Zhang et al. 2025b). Such agents could one day perform 3D reconstruction as a tool to achieve a broader goal. For autonomous driving, VLA models are being developed to incentivize reasoning and self-reflection (Yuan et al. 2025b), while physical autoregressive models are being created for robotic manipulation without action pretraining (Song et al. 2025f). Reinforcing open-vocabulary action recognition with tools is another step towards more capable and generalist agents (Yuan et al. 2025a).
Cross-Disciplinary Insights from Specialized Domains
Often, solutions developed for niche problems can inspire general-purpose algorithms. Medical Imaging, a field with unique challenges like extreme low-texture regions and noisy labels, is a fertile ground for such insights. Techniques developed for medical image segmentation, such as leveraging the frequency domain (Han et al. 2025b), estimating region uncertainty to handle noisy labels (Han et al. 2025c), and using adaptive label correction (Qian et al. 2025a), could be adapted to improve the robustness of MVS in ambiguous scenarios. Similarly, curriculum learning frameworks for imbalanced multimodal diagnosis (Qian et al. 2025b; Han et al. 2025a) could inspire new training strategies for MVS on unbalanced datasets. For cryo-electron tomography, a field with notoriously low signal-to-noise ratio, methods for self-supervised volumetric image restoration (Yang, Zhang, and Han 2021), denoising based on noise modeling and transfer (Li et al. 2022b; Yang et al. 2024d), and improved denoising with simulation-aware pretraining (Yang et al. 2024b) are highly relevant to MVS’s own struggle with noise and texturelessness. Furthermore, the use of sparse and hierarchical transformers for survival analysis on whole slide images (Yan et al. 2024a) and the fusion of multi-scale heterogeneous pathology foundation models (Yang et al. 2025) offer architectural ideas for efficient multi-scale feature aggregation. The development of evolutionary medical prompt optimization (Chen et al. 2025b) showcases a powerful method for automatically tailoring inputs to large models, a technique detailed in related publications (Chen et al. 2025c) that could be applied to guide future 3D generative systems.
Insights also come from other industrial applications. For example, efficient hybrid-supervised symmetric stereo matching networks (Zhang et al. 2023a) and effective classification methods for steel surface defects based on salient local features (Hao et al. 2024) provide battle-tested solutions for high-precision industrial vision tasks. Even a cloud framework designed for autonomous driving path planning (Yang et al. 2024a) points towards the future of MVS as a service in large-scale, distributed systems. By drawing inspiration from these diverse fields, the MVS community can continue to build more robust, efficient, and versatile 3D perception systems. Finally, the principles of contrastive learning, enhanced with diffusion models, are being used for robust depth estimation (Wang et al. 2024a), and harnessing diffusion priors for self-supervised depth estimation (Wang et al. 2025b) shows another promising direction for future MVS research.
In conclusion, by viewing MSP-MVS through the lens of broader AI advancements, we can appreciate its role not merely as an algorithmic improvement, but as a harbinger of future 3D perception systems. Its use of a foundation model for semantic guidance is indicative of a larger trend towards language-guided, context-aware intelligence. The ongoing revolutions in advanced learning paradigms and the surprising yet potent insights from cross-disciplinary fields like medical imaging provide a rich tapestry of ideas for the next generation of MVS. The future of 3D reconstruction will likely involve a deeper symbiosis with large models, more sophisticated data-centric strategies, and the integration of reasoning and action capabilities, transforming MVS from a passive reconstruction tool into an active perception module within larger intelligent agents. Our work represents a concrete step on this exciting path.